Please help. I need to find the surface area and volume of a sphere. I have a drawing of a sphere and 34in is written in the middle. Is this the radius or the diameter, can someone help me, I have been working on this problem for weeks.
Answer by fractalier(6550)

For a sphere,
V = (4/3)(pi)r^3
and
A = 4(pi)r^2
I cannot tell what the 34 in represents since I cannot see the diagram.
I am assuming it is the diameter since you say it is written in the middle of the circle. Thus the radius is half of that or 17 in. Now plug in...
V = (4/3)(pi)(17^3) = 20579.5 cubic in
A = 4(pi)(17^2) = 3631.7 sq in